Suppose Acap Corporation will pay a dividend of $2.87 per share at the end of this year and $3.05 per share next year. You expect​ Acap's stock price to be $51.44 in two years. Assume that​ Acap's equity cost of capital is 10.8%.

a. What price would you be willing to pay for a share of Acap stock​ today, if you planned to hold the stock for two​ years? (Round to the nearest​ cent.)

b. Suppose instead you plan to hold the stock for one year. For what price would you expect to be able to sell a share of Acap stock in one​ year? (Round to the nearest​ cent.)

c. Given your answer in ​(b​), what price would you be willing to pay for a share of Acap stock today if you planned to hold the stock for one​ year? How does this compare to your answer in ​(a​)? (Round to the nearest​ cent.)
Answer #1

Value of stock is equal to the present value of all future receivables

a.Price to be paid today = 2.87/(1.108) + 3.05/(1.108)2 + 51.44/(1.108)2

= $46.98

b.Price after one year = (3.05+51.44)/1.108

= $49.18

c.Price = (2.87+49.18)/1.108

= $46.98

This price is same as calculated in a part